OpenStudy (anonymous):

solve for k: f=g-h+k/7

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

no

HanAkoSolo (jamierox4ev3r):

\[f=g-h+\frac{ k }{ 7 }\] ^ Is this what your problem looks like?

OpenStudy (skullpatrol):

use the idea of "balance"

OpenStudy (anonymous):

7 is divide by all three

OpenStudy (mathstudent55):

\(f=\dfrac{g-h+k}{7} \)

OpenStudy (skullpatrol):

what you do to one side of the equation you must do the same to the other to keep it balanced

OpenStudy (anonymous):

yes

HanAkoSolo (jamierox4ev3r):

ah okay

OpenStudy (anonymous):

i know skull but im lost on the first part

HanAkoSolo (jamierox4ev3r):

remember, what is the inverse of division ^ kind of feeds into the whole "balance" idea

OpenStudy (anonymous):

so g/7 -h/7+k/7

OpenStudy (anonymous):

the change it around

OpenStudy (mathstudent55):

You can't do anything to g, h or k because they are all being divided by 7. The first step is to get rid of the division by 7 on the right side. The opposite operation to division is multiplication, so multiply both sides by 7.

HanAkoSolo (jamierox4ev3r):

basically, we're asking you what you need to multiply by in order to get rid of the denominator

HanAkoSolo (jamierox4ev3r):

which you do by multiplying both sides by 7. lel you beat me @mathstudent55 ;-;

OpenStudy (anonymous):

ok so 7*f=g-h+k

OpenStudy (skullpatrol):

Yes^

OpenStudy (mathstudent55):

\(7 \times f=7 \times \left( \dfrac{g-h+k}{7} \right) \) \(7f = g - h + k\)

OpenStudy (mathstudent55):

Good.

OpenStudy (anonymous):

then -g+h-k

OpenStudy (skullpatrol):

now subtract g from both sides

OpenStudy (anonymous):

7*f/h-g-k

OpenStudy (mathstudent55):

Now look at the right side. Anything that is being added to k, subtract from both sides. Anything that is being subtracted from k, add to both sides.

OpenStudy (anonymous):

yeah i got that part thank you

OpenStudy (mathstudent55):

\(7f = g - h + k\) \(7f = k + g - h\)

OpenStudy (fibonaccichick666):

If I may add, you know how you have the order of operations? We shorten it to PEMDAS. In order to solve for something, we reverse it and use SadMep :( <--Gotta have the frowny face :)

OpenStudy (mathstudent55):

\(7f \color{red}{-g + h}= k + g - h \color{red}{-g + h}\)

OpenStudy (skullpatrol):